40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 2401     Entry time: Fri Dec 11 17:36:37 2009
Author: kiwamu 
Type: Update 
Category: General 
Subject: IFO restoring plan 

Alberto, Jenne, Kiwamu

 

We together will lead the IFO restoring and the following is our plan.

- - - - -|- - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - -

#.0     |  measuring the free swinging spectra                     (weekend by kiwamu)   DONE

#.1     |  turn ON the PZTs for steering mirror and so on.         (Dec.14 Mon.) DONE

#. 1            |    lock around PSL  DONE

#.2     |  deal with mechanical shutter                            (Dec.14 Mon.)DONE

#.3     |  lock MCs                                                (Dec.14 Mon.)DONE

#.4     |  align the IFO                                           (Dec.15 Tue.)DONE

#.5     |  lock full IFO                                           (Dec.15 Tue.)DONE

- - - - -|- - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - -

 

Thank you.

ELOG V3.1.3-